OpenStudy (anonymous):

find the product 1.5m^6(-2m^2)^4

OpenStudy (anonymous):

1.5m^6[(-2)^4*(m^2)^4] 1.5m^6[16*m^8] 24m^14

OpenStudy (anonymous):

thank you :)

OpenStudy (anonymous):

1.5m^6 x (-2m^2)^4 = 1.5m^6 x (-2)^4 x (m^2)^4 = 1.5m^6 x 16 x m^8 = 24m^14

OpenStudy (anonymous):

welcum

OpenStudy (anonymous):

u could have atleast given me a badge:(

OpenStudy (anonymous):

hey thanks u take one

OpenStudy (anonymous):

sorry kind of new to this. i dont know what a badge is, still trying to figure out how to use this. :/